# Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting

### v1

@@ Abstract.
Spatio-temporal forecasting aims to predict the future state of dynamic systems and plays an important role in multiple fields. However, existing solutions only focus on modeling using a limited number of observed target variables. In real-world scenarios, exogenous variables can be integrated into the model as additional input features and associated with the target signal to promote forecast accuracy. Although promising, this still encounters two challenges: the inconsistent effects of different exogenous variables to the target system, and the imbalance effects between historical variables and future variables. To address these challenges, this paper introduces ExoST, a novel framework for modeling exogenous variables in spatio-temporal forecasting, which follows a “select, then balance” paradigm. Specifically, we first construct a latent space gated expert module, where fused exogenous information is projected into a latent space to dynamically select and recompose salient signals via specialized sub-experts. Furthermore, we design a siamese network architecture in which recomposed representations of past and future exogenous variables are fed into dual-branch spatio-temporal backbones to capture dynamic patterns. The outputs are integrated through a context-aware weighting mechanism to achieve dynamic balance during the modeling process. Extensive experiments on real-world datasets demonstrate the effectiveness, generality, robustness, and efficiency of our proposed framework.
1. Introduction
Spatio-temporal forecasting aims to model the dynamic evolution of spatial entities over time by leveraging historical spatio-temporal sequences to predict future states over multiple time steps (Fang et al., 2025; Shao et al., 2023; Chen and Liang, 2025a, b). This task finds critical applications in domains such as traffic management (Xia et al., 2025; Avila and Mezić, 2020), environmental monitoring (Liang et al., 2023; Wang et al., 2020), and weather forecasting (Yang et al., 2024; Bi et al., 2023). In recent years, researchers have developed various spatio-temporal (ST) neural encoders (Jin et al., 2023b, a) by combining building blocks with different relational inductive biases, such as graph convolution operator (Ma et al., 2025; Yu et al., 2017; Li et al., 2017) for spatial dependencies and sequence operator (Wu et al., 2019; Guo et al., 2019) for temporal patterns. However, these approaches predominantly rely on modeling limited observations of target system variables, focusing mainly on encoding the spatial and temporal dynamics of target signals, while overlooking a widely available yet underutilized source of information in real-world forecasting systems: spatio-temporal exogenous variables.
Spatio-temporal exogenous variables refer to contextual information that is related to the prediction system but originates from external dynamic systems (Berkani et al., 2023). Such variables may include historical information that co-occur with the target system as well as prior information about future time points provided by external auxiliary forecasting systems. These factors serve as critical complementary signals in prediction tasks, helping to reduce uncertainty while offering interpretability (Huang et al., 2025; Das et al., 2023; Wang et al., 2024; Arango et al., 2025; Olivares et al., 2023; Giganti et al., 2024). As illustrated in Fig. 1 (a), consider the sensor network of air quality monitoring stations in New York City. For any given monitoring station, air quality forecasting typically focuses on modeling various indicators of pollutant emissions (e.g., NO2, CO2) to predict future air conditions. However, air quality at a particular location may also be significantly influenced by surrounding contextual information. For instance, historical traffic congestion data in the region (e.g., traffic flow) could lead to sudden spikes in pollutant levels, while stable weather conditions predicted for the future (e.g., wind speed) may indicate a risk of continued deterioration in air quality.
Despite the considerable potential of incorporating exogenous information, two major challenges persist in practice:
✦ Challenge I: Inconsistent Variable Effects: In real-world scenarios, exogenous variables often exert varying degrees of influence on the prediction target. As shown in Figure 1(b), different exogenous variables display distinct temporal dynamics and correlations with the target signal. Some variables (e.g., traffic flow) may have substantial predictive value, while others (e.g., traffic incidents) may be marginally relevant or even introduce noise. Fig. 2 (a) further illustrates this phenomenon by showing the Pearson correlation coefficients between various exogenous variables and the target variable across randomly selected stations in the AQI-19 dataset. Despite this variability, existing approaches often adopt a one-size-fits-all strategy (Arango et al., 2025; Giganti et al., 2024), directly concatenating exogenous variables with input features, without selectively modeling their individual contributions. This lack of mechanism-aware integration limits performance and can even degrade predictive accuracy.
✦ Challenge II: Imbalanced Type Effects: In addition to inter-variable disparities, there also exist imbalanced type differences between past and future exogenous variables from different temporal contexts. As shown in Fig. 1(c), exogenous information may stem from historically perceived signals (e.g., past traffic data) or from proactively forecasted future signals (e.g., weather predictions), leading to temporal distribution asymmetry and context-dependent impacts on prediction outcomes. Fig. 2 (b) clearly demonstrates the distinct distribution patterns of normalized values across different types of exogenous variables in the AQI-19 dataset. Existing methods typically adopt joint modeling strategies that fail to distinguish the distinct influences of these two types of exogenous variables on the target system (Wang et al., 2024; Das et al., 2023; Olivares et al., 2023; Zhou et al., 2025). Moreover, even under decoupled modeling settings, simple weighted fusion is insufficient to effectively balance their heterogeneous effects.
To address the aforementioned challenges, we propose a novel exogenous variable-enhanced spatio-temporal forecasting framework, ExoST, which is built upon a core “select, then balance” paradigam to systematically exploit and integrate the contributions of exogenous variables, thereby improving overall predictive performance. The framework adopts a modular and flexible design that allows seamless integration with a wide range of mainstream spatio-temporal backbone networks. Specifically, ❶ we introduce an adaptive selection mechanism through a latent-space gated expert module, which projects both endogenous and multi-source exogenous variables into a condition latent space and activates specialized expert sub-networks via a gating mechanism. This enables the model to automatically select and combine the most informative exogenous factors. The module exhibits strong representational capacity and adaptability, effectively capturing relevant signals while suppressing irrelevant or redundant information, thus alleviating inconsistencies caused by variable heterogeneity. To address the imbalance between types of exogenous variables, ❷ we design a dual-branch siamese network architecture that independently models past and future exogenous information. These recomposed representations are fed into dual versions of any given spatio-temporal backbone to learn evolutionary patterns, and a context-aware balancer is employed to dynamic fusion and produce the final prediction.
@@ 1. Introduction
Adaptive Selection Mechanism for Inconsistent Variable Effects: We propose a Latent Space Gated Expert Module that effectively handle the latent associations among multi-source exogenous variables. Through an expert selection mechanism, it identifies and emphasizes the most predictive external signals, enhancing the quality of the semantic representation input.
Balance Modeling Strategy for Imbalanced Type Effects: We design a Siamese network architecture that separately models past and future exogenous information branch, enabling dynamic prediction balancing via a context-aware fusion gating mechanism.
A Flexible Plug-and-Play Framework: ExoST provides a general and flexible enhancement framework for spatio-temporal forecasting. It can be broadly applied to various backbones and consistently outperforms existing methods across multiple benchmarks, demonstrating both practical value and scalability.
2. Preliminaries
2.1. Related Work
2.1.1. Spatio-Temporal Forecasting.
@@ 2.1. Related Work
Predicting future spatio-temporal signals from historical observations, a spatially extension of multivariate time series forecasting, has evolved significantly. Early statistical (Box and Pierce, 1970; Biller and Nelson, 2003) or independent time series models (Nie et al., 2023; Wu et al., 2022; Qiu et al., 2025; Seo and Lim, 2025; Fan et al., 2025) at each location proved insufficient for complex spatio-temporal dynamics (Shao et al., 2024), which involve intricate spatial correlations and heterogeneity (Chen and Liang, 2025a, b). This led to hybrid architectures (Cao et al., 2025; Kong et al., 2024; Ma et al., 2025; Liu et al., 2023; Shao et al., 2022; Wu et al., 2019; Guo et al., 2019; Li et al., 2017) that integrate spatial (e.g., graph operators for node interactions) and temporal (e.g., sequence operators for long-term trends) modeling to jointly learn rich spatio-temporal representations for accurate prediction. However, a critical limitation persists: these models largely focus on the system’s internal dynamics while neglecting exogenous variables—external inputs that are independent of the system’s state but directly influence its evolution, and whose inclusion is vital for robust and accurate predictions.
2.1.2. Exogenous Variables Modeling.
In practical time series forecasting scenarios, exogenous information is already integrated into the model as additional input features and concatenated with the target signal, enabling the model to learn the correlation between external factors and the target variable. Earlier studies explored statistical methods such as ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) that can understand the relationship between exogenous and endogenous sequences and autoregression. Advanced deep learning models discuss how to use future values of exogenous variables (Das et al., 2023; Olivares et al., 2023) and how to adapt to time lags and missing exogenous variable data scenarios (Wang et al., 2024; Zhou et al., 2025). In addition, some works (Huang et al., 2025; Liu et al., 2024; Jin et al., [n. d.]) also discuss how to use language models to align the semantic information of exogenous variables for enhancement. Limited works have explored the integration of exogenous variables in spatio-temporal forecasting scenarios, but they only focus on specific task scenarios and model structure design (Giganti et al., 2024; Dong et al., 2023; Zagouras et al., 2015; Lindström et al., 2014). In contrast, we systematically explore how to effectively model exogenous variables in spatio-temporal forecasting scenarios and adaptively weigh the impact of different exogenous variables.
2.2. Problem Formulation
Definition 1.
(Spatio-Temporal Data). Spatio-temporal data refers to observational records that capture how dynamic systems evolve in space and time. Each data point is associated with a specific geographic location and a specific timestamp. This type of data can usually be represented as a temporal sequence , where each is a snapshot of the spatial observation at time . The spatial is a sensor network consisting of discrete nodes, and each node has features at time (e.g., pollutant indicators CO2, NO2, etc. from a dynamic air monitoring systems.).
Definition 2.
(Exogenous Variables). Exogenous variables are external system factors that are independent of the spatio-temporal forecasting and are free of the internal dynamics. During the modeling process, these variables are fed into the model as known inputs (whether actual observations or predicted values). Typically, these data hold the same structure as spatio-temporal data, denoted as , and can be further subdivided into past exogenous variables (e.g., traffic flow and traffic speed at past moments) and future exogenous variables (e.g., future temperature and precipitation forecasts) according to their time attributes.
Problem 0.
@@ 2.2. Problem Formulation
(ST Forecasting with Exogenous Variables). Our goal is to learn an optimal prediction model for predicting future spatio-temporal signals from historical observations, optionally using exogenous inputs (and an optional graph structure prior ). Let be a training dataset sampled from a spatio-temporal process, where represents the features of nodes for the past time steps, and represents the target for the future time steps. Traditional spatio-temporal prediction optimizes the model parameters by minimizing the expected prediction loss:
where is the prediction model and measures the prediction error. When exogenous input is available, the goal becomes:
where and denote the exogenous inputs in the past and future, respectively. and are the corresponding feature dimensions.
3. Methodology
3.1. Overview
@@
The ExoST is built on the “select-then-balance” paradigm, as shown in Figure 3. The process consists of two consecutive stages, which are used to address the two core challenges in exogenous-augmented spatio-temporal forecasting mentioned above: (i) inconsistent variable effects and (ii) unbalanced type effects.
Select stage (§3.2). A lightweight Latent-Space Gated Expert module projects heterogeneous exogenous variables into a condition latent manifold, and then activates a Mixture-of-Experts Selector to dynamically re-weight and re-compose the most meaningful signals. The output is two refined, exogenous-aware representations and , one for each type.
Balance stage (§3.3). A Siamese dual-branch architecture processes the two representations in parallel through the existing spatio-temporal backbones, and a Context-Aware Balancer adaptively balances their contributions based on instantaneous context. The final forecast is produced by a read-out layer on top of the fused hidden state.
Both stages are fully differentiable and backbone-agnostic: any modern ST encoder can be plugged in without architectural change.
3.2. Select: Latent-Space Gated Expert Module
@@
The selection stage aims to address inconsistent variable effects by extract compact, adaptive semantic representations from input endogenous variables and heterogeneous, multi-source exogenous variables, laying the foundation for the subsequent balancing stage. This requires a carefully designed two-step process: (i) semantic alignment of diverse exogenous and endogenous variables in a unified semantic space via conditional embeddings; (ii) adaptive selection through a mixture-of-experts selector that reweights and combines existing features to mitigate antagonism caused by inconsistent variable effects.
3.2.1. Conditional Embedding
@@
Conditional embeddings specifically address the semantic space alignment of endogenous variables with two types of exogenous variables. For each exogenous type (i.e., past, future), we first project the raw exogenous tensor and the target history input into a condition latent space via:
where is implemented as affine-transformation block:
where , are learned projections that align heterogeneous feature dimensions into a latent semantic space. Specifically, consistent with common settings in existing literature (Shao et al., 2024; Jin et al., 2024), the lengths of past and future steps are typically uniform. Therefore, to maintain this setting, we perform a direct element-wise add fusion. For more arbitrary cases, zero-padding preprocessing trick can also easily ensure alignment. Subsequently, non-linear activation functions project endogenous variables into latent conditional semantic spaces, derived from different types of exogenous variables, yielding conditional representations and , respectively.
3.2.2. Mixture-of-Experts Selector
As discussed, inconsistent variable effects between exogenous and endogenous variables can induce antagonistic or synergistic semantic degeneration. This manifests as channel entanglement in the derived conditional representations, where semantic mixing may lead to rank-deficient representations, complicating the learning process. To address this, we attach a mixture-of-experts gating (Jacobs et al., 1991; Shazeer et al., 2017; Fedus et al., 2022) module on top of each . Concretely, we maintain expert projections and a gating network parameterized by a single linear layer followed by function. For each conditional representation input, the gating score is applied at different time steps and locations and is calculated as follows:
@@ 3.2.2. Mixture-of-Experts Selector
where is projection layer projection layer, along with a normalized activation function, calculates each expert’s weight. This controls the weighting of each expert for the current input, simulating various semantic combinations. The final representation is then obtained through weighted combination:
This strategy explicitly disentangles representations using a latent bottleneck and selective recombination, rather than forcing the model to learn entangled mappings. This latent space gating method allows us to model multiple interactions between exogenous variables and the target representation within the fused conditional space, thereby adaptively acquiring suitable representation input and for the backbone network.
3.3. Balance: Dual-Branch Adaptive Fusion
The balancing stage aims to address the unbalanced type effect, that is, how to properly process the two semantic representations in the previous stage. This requires appropriately learning spatio-temporal dynamic patterns and effectively fusing them to achieve the final prediction. However, a key challenge arises from the inherent difference in temporal distribution between past and future exogenous information, complicating their effective fusion. To address this, we designed two crucial steps: (i) independent extraction of dynamic spatio-temporal features for both inputs via parallel siamese network branches. (ii) adaptive combination through a context-aware fusion gating mechanism that generates dynamic fusion weights to mitigate the imbalanced type effects.
3.3.1. Siamese ST Encoders
To effectively process the distinct characteristics of past- and future-conditioned semantic representations, we adopt a siamese architecture. This involves two identical spatio-temporal (ST) backbone encoders. Each encoder , which can be instantiated with any existing spatio-temporal neural network (Bai et al., 2020; Guo et al., 2019; Wu et al., 2019; Li et al., 2017; Yu et al., 2017; Satorras et al., 2022; Gao and Ribeiro, 2022), is tasked with capturing the complex spatial correlations and temporal dependencies within its respective input stream. The parallel processing yields two specialized representations:

### v2

# Select, then Balance: Exploring Exogenous Variable Modeling of Spatio-Temporal Forecasting

@@ Abstract
Spatio-temporal (ST) forecasting is critical for dynamic systems, yet existing methods predominantly rely on modeling a limited set of observed target variables. In this paper, we present the first systematic exploration of exogenous variable modeling for ST forecasting, a topic long overlooked in this field. We identify two core challenges in integrating exogenous variables: the inconsistent effects of distinct variables on the target system and the imbalance effects between historical and future data. To address these, we propose ExoST, a simple yet effective exogenous variable modeling general framework highly compatible with existing ST backbones that follows a “select, then balance” paradigm. Specifically, we design a latent space gated expert module to dynamically select and recompose salient signals from fused exogenous information. Furthermore, a siamese dual-branch backbone architecture captures dynamic patterns from the recomposed past and future representations, integrating them via a context-aware weighting mechanism to ensure dynamic balance. Extensive experiments on real-world datasets demonstrate the ExoST ’s effectiveness, universality, robustness, and efficiency.
onedeanxxx@gmail.com, ∗yuxliang@outlook.com
1 Introduction
Spatio-temporal forecasting aims to learn the evolution of structured systems whose states vary across both space and time (shao2024exploring; chen2025eac; chen2025learning). This task is fundamental to traffic management (avila2020data), environmental monitoring (liang2023airformer), and meteorology (bi2023accurate). Recent advances have focused on spatio-temporal (ST) encoders (jin2023spatio; jin2024survey) that integrate graph neural operators (yu2017spatio; ma2025less) for learning spatial dependencies and sequence operators (wu2019graph; guo2019attention; gao2023spatio; han2024bigst) for modeling temporal patterns. However, existing methods predominantly focus on the dynamics of target variables, largely neglecting a critical source of auxiliary information: spatio-temporal exogenous variables.
Spatio-temporal exogenous variables encompass contextual data from external dynamic systems (berkani2023spatio), including historical covariates and auxiliary future forecasts. These signals complement target predictions by reducing uncertainty and enhancing interpretability (olivares2023neural; wang2024timexer; huang2025exploiting; Arango2025ChronosX). As illustrated in Figure 1 (a), air quality forecasting relies not only on historical pollutant indicators (e.g., NO2, CO2) but also on the environmental context. For instance, historical traffic congestion may explain sudden emission spikes, while future weather forecasts (e.g., wind speed) serve as precursors for pollutant dispersion or accumulation.
Despite the promise of exogenous information, their systematic utilization in ST forecasting remains surprisingly overlooked. We attribute this gap to two primary challenges:
Challenge I: Inconsistent Variable Effects: Exogenous variables exhibit heterogeneous correlations with prediction targets. As shown in Figure 1 (b) (Pearson correlation coefficients between exogenous variables and the target variable across randomly stations in AQI-19), variables vary from highly predictive (e.g., traffic flow) to irrelevant or noisy (e.g., traffic incidents). Existing approaches typically employ a uniform concatenation strategy (giganti2024back; Arango2025ChronosX), which fails to distinguish informative signals from noise. This lack of selective modeling limits representation capability and degrades predictive accuracy.
Challenge II: Imbalanced Type Effects: In addition to inter-variable disparities, distinct distributional asymmetry exists between historical signals and proactive future forecasts. Figure 1 (c) highlight the disparate distribution patterns of these data types in AQI-19. Current methods often rely on joint modeling or simple weighted fusion (das2023long; olivares2023neural; wang2024timexer; zhou2025crosslinear), which fail to effectively decouple and balance the different impacts of past and future contexts on the target system.
To address these limitations, we propose ExoST, a novel framework governed by a “select-then-balance” paradigm designed to systematically exploit exogenous variables. ExoST features a modular design compatible with diverse spatio-temporal backbones. We introduce a Latent Space Gated Expert Module that projects endogenous and exogenous variables into a shared latent space, utilizing a gating mechanism to adaptively select informative factors while suppressing noise. To mitigate type imbalance, we design a dual-branch Siamese architecture that decouples the modeling of past and future exogenous contexts. These representations are processed by dual backbones and integrated via a context-aware balancer to generate the final prediction.
@@ 1 Introduction
Problem Level: We present the first systematic formalization and exploration of exogenous variable modeling for ST forecasting, addressing a long-neglected gap therein.
Technical Level: We identify the key challenges in modeling exogenous variables and propose ExoST, a simple yet effective general framework with targeted modules. ExoST demonstrates high compatibility with existing spatio-temporal backbones without any modification.
Evaluation Level: We systematically evaluate ExoST across multiple aspects, verifying its efficacy, universality, robustness, design rationality, efficiency, and lightweight features.
2 Related Work
@@
Spatio-Temporal Forecasting. Predicting future spatio-temporal signals from historical observations, a spatially extension of multivariate time series forecasting, has evolved significantly. Early statistical (box1970distribution; biller2003modeling) or independent time series models (nie2023time; wu2023timesnet; qiu2025duet) at each location proved insufficient for complex spatio-temporal dynamics (shao2024exploring), which involve intricate spatial correlations and heterogeneity (chen2025eac; chen2025learning). This led to hybrid architectures (li2017diffusion; wu2019graph; liu2023spatio; chen2025information; cao2025spatiotemporal; ma2025robust) that integrate spatial (e.g., graph operators for node interactions) and temporal (e.g., sequence operators for long-term trends) modeling to jointly learn rich spatio-temporal representations for accurate prediction. However, a critical limitation persists: these models largely focus on the system’s internal dynamics while neglecting exogenous variables, external inputs that are independent of the system’s state but directly influence its evolution, and whose inclusion is vital for robust and accurate predictions.
Exogenous Variables Modeling. Distinct from our ST forecasting setting, which models multiple systems as exogenous variables, existing literature primarily focuses on simple covariate modeling within multivariate time series forecasting to capture correlations between observations and target variables. Early statistical methods, such as ARIMAX (williams2001multivariate) and SARIMAX (vagropoulos2016comparison), addressed the relationships among exogenous series, endogenous series, and autoregressive components. More recent deep learning approaches have investigated utilizing future exogenous values (das2023long; olivares2023neural), handling time lags and missing data (wang2024timexer; zhou2025crosslinear), and leveraging language models to align semantic information for representation enhancement (liu2024spatial; huang2025exploiting). While limited research integrates exogenous variables into ST forecasting, as summarized in Table 1, these efforts are restricted to specific ST tasks and model architectural designs (lindstrom2014flexible; zagouras2015role; dong2023spatiotemporal; giganti2024back; ma2025causal; ma2025spatiotemporal; wang2025unifying). In contrast, we provide the first systematic formalization and exploration of modeling exogenous variables within this context, proposing a simple yet effective general framework.
| Category | Example Works | Typical Integration | Universality | Simplicity | Robustness | Efficiency |
| Statistical Models | ARIMAX (williams2001multivariate), SARIMAX (vagropoulos2016comparison) | Linear Regression | ✗ (Strict Assump.) | ✓ | ✗ | |
| Specific TS Models | TiDE (das2023long), TimeXer (wang2024timexer) | Concatenation / Cross-Attn | ✗ (TS specific) | ✗ | ✓ | |
| LLM-based Models | ST-LLM (liu2024spatial), ExoLLM (huang2025exploiting) | Semantic Embedding | ✓ (Semantic Bridge) | ✗ | ✗ | |
| Specific ST Models | SEEformer (dong2023spatiotemporal), MAGCRN (giganti2024back) | Task-Specific Fusion | ✗ (ST specific) | ✗ | ✓ | |
| General ST Framework | ExoST (Ours) | Select, then Balance | ✓ (Any ST Backbone) | ✓ | ✓ | ✓ |
3 Preliminaries
Definition 1. (Spatio-Temporal Data). Spatio-temporal data records the evolution of dynamic systems across space and time. Formally, we represent this as a sequence , where denotes the system state at time . The spatial domain is modeled as a network with nodes, where each node observes features (e.g., pollutant indicators such as NO2, etc. in air quality system).
Definition 2. (Exogenous Variables). Exogenous variables are external factors independent of the target system’s internal dynamics, serving as auxiliary inputs. Denoted as , these variables share the structural format of ST data. comprises past exogenous (e.g., historical traffic flow) and future exogenous variables (e.g., weather forecasts), distinguished by their temporal availability.
@@ 3 Preliminaries
Problem. (ST Forecasting with Exogenous Variables). We aim to predict future spatio-temporal signals given historical observations, optional exogenous inputs, and a graph prior . Let denote the training set, where and represent historical features (length ) and prediction targets (length ), respectively. Standard forecasting optimizes parameters by minimizing the expected loss:
where is the model and denotes the error metric. With exogenous inputs , the objective changes to:
where and correspond to past and future exogenous features with dimensions and . In practice, according to (jin2023spatio; jin2024survey; shao2024exploring), the feature to be predicted is usually only the target variable, i.e., in is equal to 1.
4 Methodology
@@
The ExoST is built on the “select-then-balance” paradigm, as shown in Figure 2. The process consists of two consecutive stages, which are used to address the two core challenges in exogenous-aware ST forecasting mentioned above: (i) inconsistent variable effects and (ii) unbalanced type effects. Both stages are fully differentiable and backbone-agnostic: any modern ST encoder can be plugged in without architectural change.
4.1 Select: Latent-Space Gated Expert
@@
The selection stage aims to address inconsistent variable effects by extract compact, adaptive semantic representations from input endogenous variables and heterogeneous, multi-source exogenous variables, laying the foundation for the subsequent balancing stage. It operates in two key stages.
@@
Conditional Embedding. Conditional embeddings specifically address the semantic space and dimension size alignment of endogenous variables with two types of exogenous variables. For each type (i.e., past, future), we project the raw exogenous tensor and historical input into a condition latent space:
where acts as an affine transformation block:
with learnable projections and . Following standard protocols (shao2024exploring; jin2024survey), we assume uniform sequence lengths of past and future steps to enable direct element-wise addition. For more arbitrary cases, the zero-padding preprocessing trick can also easily ensure alignment. This process yields the conditional representations .
Gated Expert Selector. To mitigate semantic degeneration and channel entanglement arising from inconsistent variable effects, we incorporate a Gated Expert Selector module for each . Specifically, we employ expert projections and a gating network . The gating scores , which determine the contribution of each expert based on specific spatio-temporal contexts, are computed via:
@@
where is a learnable projection matrix, along with a normalized activation function, calculates each expert’s weight. The refined representation is subsequently synthesized through a weighted aggregation:
This mechanism explicitly disentangles representations via a latent bottleneck and selective recombination, adaptively deriving robust inputs and for the backbone rather than relying on entangled mappings.
Analysis. Unlike mixture-of-experts (MoEs) in modern language models (shazeer2017outrageously; xue2024openmoe) designed for efficient sparse parallelism, ours resembles early recommender-system MoEs (ma2018modeling) that capture diverse feature patterns (e.g., rain–rush-hour or fog–low-flow). It activates specific experts under local exogenous conditions, resolving inconsistency by decoupling conflicting signals and enabling specialized processing paths (More formal analysis in Appendix A.1). Fig. 4 shows the gating probability distributions across time steps in the AQI-19 experiment, highlighting the smooth composition of expert allocation. Empirically, we also validate the module’s efficacy across diverse scenarios in Section. 5.5.
4.2 Balance: Dual-Branch Adaptive Fusion
The balancing stage addresses unbalanced type effects arising from the distributional discrepancies between past and future exogenous contexts. To effectively fuse these heterogeneous signals, we employ parallel Siamese networks for independent ST feature extraction, coupled with a context-aware balance mechanism for adaptive fusion.
Siamese ST Encoders. To process the distinct characteristics of past- and future-conditioned semantic representations, we employ a Siamese architecture comprising parallel encoders. Each encoder , adaptable to arbitrary spatio-temporal backbones (yu2017spatio; li2017diffusion; guo2019attention; wu2019graph; Bai2020Adaptive; Gao2022On; Satorras2022Multivariate), captures intrinsic spatial correlations and temporal dependencies via: